# Primes and absolutely or non-absolutely irreducible elements in atomic domains

### v1

@@ Abstract.
V. Fadinger was supported by the Austrian Science Fund (FWF): W1230
S. Frisch and D. Windisch were supported by the Austrian Science Fund (FWF): P 30934
D. Smertnig was supported by the Slovenian Research and Innovation Agency (ARIS): Grant P1-0288 and the Austrian Science Fund (FWF): P 36742.
@@ 1. Introduction
Monoid examples for all eight scenarios are easy to find, compared to examples in integral domains. For instance, Baginski and Kravitz [7] provide examples of non-factorial monoids whose irreducibles are all absolutely irreducible.
We will show that all eight scenarios occur in atomic domains. Some cases are trivial: Atomic domains without any irreducible elements, let alone absolutely irreducible or prime elements, are just fields. (There are also non-atomic domains without irreducible elements, the so-called antimatter domains [2].)
Atomic domains with primes and no other irreducible elements are just UFDs (that are not fields), as mentioned. For the remaining six non-trivial combinations we now proceed to give examples.
@@
Section Example Irreducible but not Absolutely irreducible Prime
@@
Section 6 $\mathbb{Z}[\sqrt{-14}]$ + + +
Section 4 $\operatorname{Int}(\mathcal{O}_{K})$ + + -
Section 5 $\mathbb{R}+X\mathbb{C}[X]$ + - +
Section 3 $\mathbb{R}+X\mathbb{C}[[X]]$ + - -
Section 8 $R_{2}$ + +
Section 7 $R_{1}$ + -
Section 1 UFDs - - +
Section 1 Fields - - -
@@ 2.1. Factorization terms
In this section, we recall some concepts related to factorization. For a comprehensive introduction to non-unique factorizations, we refer to the textbook by Geroldinger and Halter-Koch [15].
@@
Let $(H,\cdot)$ be a cancellative commutative monoid.
@@
  2. (ii)
    A factorization of $r\in H$ is an expression
    $r=a_{1}\cdots a_{n},$
@@
  3. (iii)
@@
    $r=a_{1}\cdots a_{n}=b_{1}\cdots b_{m},$ (1)
@@
  7. (vii)
    $(H,\cdot)$ is half-factorial if $H$ is atomic and any two factorizations of an element have the same number of irreducible factors.
@@
    An irreducible element $r\in\leavevmode\nobreak\ H$ is called absolutely irreducible (or, a strong atom), if for all natural numbers $n$, every factorization of $r^{n}$ is essentially the same as $r^{n}=r\cdots r$.
@@ 2.2. Transfer homomorphisms
Transfer homomorphisms are a key tool in factorization theory. They are used to study non-unique factorization in a domain (or monoid) using a simpler “model” monoid. In this section, we show that transfer homomorphism preserve absolute irreducibility (in the forward direction).
@@
[15, Definition 3.2.1] Let $H$ and $M$ be commutative monoids. A homomorphism $\theta\colon H\longrightarrow M$ is called a transfer homomorphism if it has the following properties:
@@
    If $u\in H$ and $b$$c\in M$ such that $\theta(u)=bc$, then there exists $v$$w\in H$ such that $u=vw$ and $\theta(v)\sim b$ and $\theta(w)\sim c$.
Here and elsewhere in this paper, the group of units of the monoid $H$ is denoted by $H^{\times}$.
Fact 2.4.
@@
Let $\theta\colon H\longrightarrow M$ be a transfer homomorphism, and $c\in H$ an irreducible element. If $c$ is absolutely irreducible in $H$, then $\theta(c)$ is absolutely irreducible in $M$.
@@
If $\theta(c)$ is not absolutely irreducible in $M$, then there exists an irreducible element $\theta(b)$ of $M$, not associated to $\theta(c)$, that divides $\theta(c)^{m}$ for some $m\in\mathbb{N}$. It follows that $b$ is an irreducible element of $H$, not associated to $c$, which divides $c^{m}$. ∎
@@
The converse of Lemma 2.5 does not hold. A non-absolutely irreducible element of $H$ may be mapped to an absolutely irreducible element of $M$ by a transfer homomorphism $\theta\colon H\longrightarrow M$. We illustrate this by an example: If $H$ is a commutative monoid and $M=(\mathbb{N}_{0},+)$, then the function
@@
$a\longmapsto\ell(a)$
is a transfer homomorphism.
If $c\in H$ is irreducible in $H$, then $\theta(c)$ is irreducible in $M$, by Fact 2.4. The unique irreducible of $M$ is $1$ (which is prime). Hence $\theta(c)=1$ for all irreducible $c\in H$, and in particular, every irreducible in $M$ is absolutely irreducible. However, there exist half-factorial monoids $H$ that contain non-absolutely irreducibles, for instance $H=\mathcal{O}_{K}\setminus\{0\}$ with $\mathcal{O}_{K}$ a ring of algebraic integers whose class group is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ (see Fact 6.2 below).
@@ 2.3. Krull monoids
A cancellative commutative monoid is a Krull monoid if it is $v$-noetherian and completely integrally closed. A Krull domain is an integral domain $D$ such that $D\setminus\{0\}$ is a Krull monoid. We refer to [15, Chapter 2] for the algebraic theory of Krull monoids.
Let $H$ be a Krull monoid and let $\mathfrak{X}(H)$ denote the set of nonempty divisorial prime (semigroup) ideals. The nonempty divisorial ideals of $H$ form a free abelian monoid with basis $\mathfrak{X}(H)$ with respect to the divisorial product; the divisorial fractional ideals form a free abelian group on the same basis. Explicitly, every nonempty divisorial ideal $\mathfrak{a}$ of $H$ is uniquely expressible as a divisorial product of prime ideals
@@
The set $\{\,\mathfrak{p}\in\mathfrak{X}(H):\mathsf{v}_{\mathfrak{p}}(\mathfrak{a})>0\,\}=\{\,\mathfrak{p}\in\mathfrak{X}(H):\mathfrak{p}\supseteq\mathfrak{a}\,\}$ is the support of $\mathfrak{a}$. The class group $G=\mathcal{C}(H)$ of $H$ is the quotient of the group of divisorial fractional ideals of $H$ by the subgroup of principal fractional ideals. Let $[\mathfrak{a}]\in G$ denote the class of $\mathfrak{a}$. In factorization theory, the set $G_{0}=\{\,[\mathfrak{p}]:\mathfrak{p}\in\mathfrak{X}(H)\,\}$ of classes containing prime divisors is of central importance.
We have two main examples of Krull monoids in mind: the first are the Dedekind domains, which are Krull domains (in fact, they are precisely the one-dimensional Krull domains). In a Dedekind domain $D$, the map $\mathfrak{a}\mapsto\mathfrak{a}\setminus\{0\}$ is a bijection between ring ideals of $D$ and divisorial semigroup ideals of $D\setminus\{0\}$. The divisorial product is the usual product of ideals, and the class group is the usual one (the group of fractional ideals by principal fractional ideals).
@@
The irreducibles of $\mathcal{B}(G_{0})$ are the minimal zero-sum sequences: nonempty sequences whose sum is 00, but which do not contain a nonempty, proper zero-sum subsequence whose sum is 00.
@@ 3. Rings whose irreducible elements are all non-absolutely irreducible
Let $K_{1}\subseteq K_{2}$ be fields and $n\in\mathbb{N}$, and set $R=K_{1}+X^{n}K_{2}[[X]]$. Then the following hold.
@@
    $R$ has prime elements if and only if $K_{1}=K_{2}$ and $n=1$.
@@
For (i), suppose $K_{1}=K_{2}$ and $n=1$. Then $R$ is a unique factorization domain.
Conversely, suppose $K_{1}=K_{2}$ and $n>1$. Then the units of $R$ are the elements of $R$ with constant term in $K_{2}\setminus\{0\}$. It follows from Remark 3.1 and Fact 2.4 that the irreducible elements of $R$ are the polynomials of the form $r=uX^{m}$, where $u$ is a unit of $K_{2}[[X]]$ and $n\leq m\leq 2n-1$. Every irreducible of the form $r$ is not absolutely irreducible since for some $n\leq t\leq 2n-1$, with $t\neq m,$
@@
Suppose $K_{1}\neq K_{2}$. Then the units of $R$ are the elements of $R$ with constant term in $K_{1}\setminus\{0\}$. Similarly, the irreducible elements of $R$ are the polynomials of the form $uX^{m}$, where $u$ is a unit of $K_{2}[[X]]$ and $n\leq m\leq 2n-1$. Moreover, $u_{1}X^{m}\not\sim u_{2}X^{m}$ if $u_{1}$ is a unit in $K_{1}[[X]]$ and $u_{2}\in K_{2}[[X]]$ has its constant term in $K_{2}\setminus K_{1}$. It follows that each irreducible of the form $uX^{m}$ is not absolutely irreducible since
@@
is a factorization of $(uX^{m})^{2}$ essentially different from $uX^{m}\cdot uX^{m}$, where $c\in K_{2}\setminus K_{1}$.
For (ii), if $K_{1}=K_{2}$ and $n=1$, then $R=K_{2}[[X]]$ is a unique factorization domain, otherwise, the assertion follows from (i) and the fact that every prime element is absolutely irreducible. ∎
@@ 4. Rings with both absolutely and non-absolutely irreducible elements, but no primes
The fixed divisor of $f\in\operatorname{Int}(D)$, abbreviated $\operatorname{fd}(f)$, is the ideal of $D$ generated by $f(D)$. If the fixed divisor is a principal ideal, we say $\operatorname{fd}(f)=c$, by abuse of notation, for $\operatorname{fd}(f)=(c)$. A polynomial $f\in\operatorname{Int}(D)$ with $\operatorname{fd}(f)=1$ is called image-primitive. It is clear that $\operatorname{fd}(f)\cdot\operatorname{fd}(g)\supseteq\operatorname{fd}(fg)$, but note that the inclusion can be strict. In any case, all divisors in $\operatorname{Int}(D)$ of an image-primitive polynomial $f\in\operatorname{Int}(D)$ are image-primitive.
@@ 4.1. No prime elements
We conclude from Fact 4.2 that every non-constant polynomial in $\mathbb{Z}[x]$ has zeros modulo infinitely many primes.
@@
Now consider $G\in\operatorname{Int}(\mathbb{Z})$ non-constant, $G(x)=\frac{g}{d}$ with $g(x)\in\mathbb{Z}[x]$ and $d\in\mathbb{Z}$. Let $p\in\mathbb{Z}$ be prime such that $g$ has a zero modulo $p$ but the polynomial function induced by $g$ is not constant zero modulo $p$. (Such a prime $p$ exists because $g$ has a zero modulo infinitely many primes, but only finitely many primes divide the fixed divisor of $g$.)
Since $g\in\mathbb{Z}[x]$, the residue class of $g(r)$ modulo $p$ depends only on the residue class of $r$ modulo $p$. Let $r_{1},\ldots,r_{k}$ be a complete set of representatives of those residue classes modulo $p$ on which $g$ takes a non-zero value modulo $p$. Let $h(x)=\prod_{i=1}^{k}(x-r_{i})$.
Then, both $\frac{h(x)G(x)}{p}$ and $\frac{(G(x)+p)h(x)}{p}$ are integer-valued. This means that $G$ divides $\frac{(G(x)+p)h(x)G(x)}{p}$, but $G$ divides neither $G(x)+p$ (not even in $\mathbb{Q}[x]$) nor $\frac{h(x)G(x)}{p}$ (because $\frac{h(x)}{p}$ is not integer-valued). ∎
@@
Since $\mathbb{Z}$ is a d-ring, it follows by Fact 4.4 that $\mathcal{O}_{K}$, too, is a d-ring.

### v2

@@ Abstract.
V. Fadinger was supported by the Austrian Science Fund (FWF) [10.55776/W1230].
S. Frisch (the corresponding author) and D. Windisch acknowledge support by the Austrian Science Fund (FWF) [10.55776/P30934].
S. Nakato was supported in part by Africa-UniNet under the project P142_Uganda: Uganda-Austria Collaboration in Algebra and Geometry
D. Smertnig was funded in whole or in part by the Slovenian Research and Innovation Agency (ARIS): Grants P1-0288, J1-60025 and the Austrian Science Fund (FWF) [10.55776/P36742].
@@ 1. Introduction
Monoid examples for all eight scenarios are easy to find, compared to examples in integral domains. For instance, Baginski and Kravitz [7] provide examples of non-factorial monoids whose irreducibles are all absolutely irreducible, but not prime.
We will show that all eight scenarios occur in atomic domains.
Remark 1.1.
Some cases are trivial: Atomic domains without any irreducible elements, let alone absolutely irreducible or prime elements, are just fields. (There are also non-atomic domains without irreducible elements, the so-called antimatter domains [2].)
Atomic domains containing primes but no other irreducible elements are precisely those UFDs that are not fields, as mentioned.
For the six non-trivial combinations we now proceed to give examples.
@@
Result Example Irreducible but not Absolutely irreducible Prime
@@
6.2,6.4 $\mathbb{Z}[\sqrt{-14}]$ + + +
4.9 $\operatorname{Int}(\mathcal{O}_{K})$ + + -
5.1,5.2 $\mathbb{R}+X\mathbb{C}[X]$ + - +
3.2 $\mathbb{R}+X\mathbb{C}[[X]]$ + - -
9.1 $R_{2}$ - + +
8.1 $R_{1}$ - + -
1.1 UFDs - - +
1.1 Fields - - -
@@ 2.1. Factorization terms
We recall some concepts and terminology related to factorization. For a comprehensive introduction to non-unique factorizations, we refer to the textbook by Geroldinger and Halter-Koch [15].
@@
Let $(H,\cdot)$ be a commutative monoid.
@@
  3. (iii)
    A factorization of a non-unit $r\in H$ is an expression
    $r=a_{1}\cdot\ldots\cdot a_{n},$
@@
    The number $n$ of irreducible factors is called the length of the factorization.
  4. (iv)
@@
    $r=a_{1}\cdot\ldots\cdot a_{n}=b_{1}\cdot\ldots\cdot b_{m},$ (1)
@@
  8. (viii)
    $(H,\cdot)$ is half-factorial if $H$ is atomic and any two factorizations of an element have the same length, i.e., the same number of irreducible factors.
  9. (ix)
    In a half-factorial monoid, the length of a non-zero element $h$, denoted $\ell(h)$, is defined as the length of a factorization of $h$ into irreducibles.
Definition 2.2.
A commutative monoid $(H,\cdot)$ is called cancellative if $ab=ac$ implies $b=c$, for any $a,b,c\in H$.
The quotient group $\operatorname{\mathbf{q}}(H)$ of a cancellative monoid is the group defined on the set of equivalence classes of pairs $(a,b)\in H\times H$ with respect to the equivalence relation
$(a,b)\simeq(c,d)\Longleftrightarrow ad=bc,$
endowed with the multiplication
$\frac{a}{b}\cdot\frac{c}{d}=\frac{ab}{cd},$
where $\frac{a}{b}$ denotes the equivalence class of $(a,b)$.
@@
    An irreducible element $r\in\leavevmode\nobreak\ H$ is called absolutely irreducible (or, a strong atom), if for all natural numbers $n$, every factorization of $r^{n}$ is essentially the same as $r^{n}=r\cdot\ldots\cdot r$.
@@ 2.2. Transfer homomorphisms
Transfer homomorphisms are a key tool in factorization theory. They are used to study non-unique factorization in a domain (or monoid) using a simpler “model” monoid. In this section, we show that transfer homomorphisms preserve absolute irreducibility (in the forward direction).
@@
[15, Definition 3.2.1] Let $H$ and $M$ be cancellative commutative monoids. A monoid homomorphism $\theta\colon H\longrightarrow M$ is called a transfer homomorphism if it has the following properties:
@@
[…]
Fact 2.5.
@@
Let $\theta\colon H\longrightarrow M$ be a transfer homomorphism, and $c\in H$ an irreducible element. If $c$ is absolutely irreducible in $H$ then $\theta(c)$ is absolutely irreducible in $M$.
@@
If $\theta(c)$ is not absolutely irreducible in $M$, then there exists an irreducible element $a\in M$, not associated to $\theta(c)$, that divides $\theta(c)^{m}$ for some $m\in\mathbb{N}$. By the first of the defining properties of a transfer homomorphism, we may assume $a=\theta(b)$ for some $b\in H$. Since $\theta(b)$ is irreducible, it follows by the second of the defining properties of a transfer homomorphism that this $b$ is an irreducible element of $H$, dividing $c^{m}$. Also, $b$ cannot be associated to $c$, because otherwise $\theta(b)$ would be associated to $\theta(c)$. ∎
@@
The converse of Lemma 2.6 does not hold. A non-absolutely irreducible element of $H$ may be mapped to an absolutely irreducible element of $M$ by a transfer homomorphism $\theta\colon H\longrightarrow M$. We illustrate this by an example: If $H$ is a half-factorial commutative monoid and $M=(\mathbb{N}_{0},+)$, then the function
@@
$a\longmapsto\ell(a),$
where $\ell(a)$ denotes the length of $a$ (that is, the number of irreducibles in a factorization of $a$), is a transfer homomorphism.
If $c\in H$ is irreducible in $H$, then $\theta(c)$ is irreducible in $M$, by Fact 2.5. The unique irreducible of $M$ is $1$ (which is prime). Hence $\theta(c)=1$ for all irreducible $c\in H$, and in particular, every irreducible in $M$ is absolutely irreducible. However, there exist half-factorial monoids $H$ that contain non-absolutely irreducibles, for instance $H=\mathcal{O}_{K}\setminus\{0\}$ with $\mathcal{O}_{K}$ a ring of algebraic integers whose class group is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ (see Fact 6.3 below).
@@ 2.3. Krull monoids
Recall that an integral domain is Krull if and only if it is completely integrally closed and Mori. The concepts involved in this characterization depend only on the multiplicative monoid and can thus be used to define Krull monoids.
A cancellative commutative monoid $H$ is called completely integrally closed if it contains all elements of $\operatorname{\mathbf{q}}(H)$ that are almost integral over $H$, that is, those $c\in\operatorname{\mathbf{q}}(H)$ for which there exists some $d\in H$ such that for all $n\in\mathbb{N}$, $dc^{n}\in H$.
Likewise, a cancellative commutative monoid $(H,\cdot)$ is Mori if it satisfies the ascending chain condition for divisorial ideals. Here, a divisorial ideal is defined just like a divisorial ideal of an integral domain, with the quotient group $\operatorname{\mathbf{q}}(H)$ of the cancellative monoid taking the place of the quotient field of an integral domain.
A cancellative commutative monoid is a Krull monoid if it is completely integrally closed and Mori.
With this definition of Krull monoid, an integral domain $D$ is Krull if and only if $D\setminus\{0\}$ is a Krull monoid [15, Thm. 2.10.2.3].
We refer to [15, Chapter 2] for the algebraic theory of Krull monoids and for more details on the terminology we just introduced.
Let $H$ be a Krull monoid and let $\mathfrak{X}(H)$ denote the set of nonempty divisorial prime (semigroup) ideals. The nonempty divisorial ideals of $H$ form a free abelian monoid with basis $\mathfrak{X}(H)$ with respect to the divisorial product; the nonempty divisorial fractional ideals form a free abelian group on the same basis. Explicitly, every nonempty divisorial ideal $\mathfrak{a}$ of $H$ is uniquely expressible as a divisorial product of prime ideals
@@
The set $\{\,\mathfrak{p}\in\mathfrak{X}(H):\mathsf{v}_{\mathfrak{p}}(\mathfrak{a})>0\,\}=\{\,\mathfrak{p}\in\mathfrak{X}(H):\mathfrak{p}\supseteq\mathfrak{a}\,\}$ is the support of $\mathfrak{a}$. The class group $G=\mathcal{C}(H)$ of $H$ is the quotient of the group of nonempty divisorial fractional ideals of $H$ by the subgroup of principal fractional ideals. Let $\mathfrak{a}]\in G$ denote the class of $\mathfrak{a}$. In factorization theory, the set $G_{0}=\{\,[\mathfrak{p}]:\mathfrak{p}\in\mathfrak{X}(H)\,\}$ of classes containing prime divisors is of central importance.
We have two main examples of Krull monoids in mind: the first are the Dedekind domains, which are Krull domains. In fact, Dedekind domains (apart from fields, which usually count as Dedekind domains, too) are precisely the one-dimensional Krull domains [15, Thm. 2.10.6]. In a Dedekind domain $D$, the map $\mathfrak{a}\mapsto\mathfrak{a}\setminus\{0\}$ is a bijection between ring ideals of $D$ and divisorial semigroup ideals of $D\setminus\{0\}$.
The divisorial product is the usual product of ideals, and the class group is the usual one (the group of fractional ideals by principal fractional ideals).
@@
The irreducibles of $\mathcal{B}(G_{0})$ are the minimal zero-sum sequences: nonempty sequences whose sum is 0, but which do not contain a nonempty proper subsequence whose sum is 0.
@@ 3. Rings whose irreducible elements are all non-absolutely irreducible
Let $K_{1}\subseteq K_{2}$ be fields and $n\in\mathbb{N}$, and set $R=K_{1}+X^{n}K_{2}[[X]]$. Then the following are equivalent.
@@
    $R$ is a UFD.
  3. (iii)
    every irreducible of $R$ is prime.
  4. (iv)
    $R$ has a prime element.
  5. (v)
    $R$ has an absolutely irreducible element.
@@
The implications (i) $\Longrightarrow$ (ii) $\Longrightarrow$ (iii) $\Longrightarrow$ (iv) $\Longrightarrow$ (v) follow immediately.
For (v) $\Longrightarrow$ (i), suppose $K_{1}=K_{2}$ and $n>1$. Then the units of $R$ are the elements of $R$ with constant term in $K_{2}\setminus\{0\}$. It follows from Remark 3.1 and Fact 2.5 that the irreducible elements of $R$ are the polynomials of the form $r=uX^{m}$, where $u$ is a unit of $K_{2}[[X]]$ and $n\leq m\leq 2n-1$. Every irreducible of the form $r$ is not absolutely irreducible since for any $n\leq t\leq 2n-1$, with $t\neq m,$
@@
Suppose $K_{1}\neq K_{2}$. Then the units of $R$ are the elements of $R$ with constant term in $K_{1}\setminus\{0\}$. Similarly, the irreducible elements of $R$ are the polynomials of the form $uX^{m}$, where $u$ is a unit of $K_{2}[[X]]$ and $n\leq m\leq 2n-1$. Moreover, if $u_{1},u_{2}$ are units of $K_{2}[[X]]$, then $u_{1}X^{m}\not\sim u_{2}X^{m}$ if $u_{1}u^{-1}_{2}$ has a constant term in $K_{2}\setminus K_{1}$. It follows that each irreducible of the form $uX^{m}$ is not absolutely irreducible since
@@
is a factorization of $(uX^{m})^{2}$ essentially different from $uX^{m}\cdot uX^{m}$, where $c\in K_{2}\setminus K_{1}$. ∎
@@ 4. Rings with both absolutely and non-absolutely irreducible elements, but no primes
Definition 4.1.
The fixed divisor of $f\in\operatorname{Int}(D)$, abbreviated $\operatorname{fd}(f)$, is the ideal of $D$ generated by $f(D)$. If the fixed divisor is a principal ideal, we say $\operatorname{fd}(f)=c$, by abuse of notation, for $\operatorname{fd}(f)=(c)$. A polynomial $f\in\operatorname{Int}(D)$ with $\operatorname{fd}(f)=1$ is called image-primitive.
It is clear that $\operatorname{fd}(f)\cdot\operatorname{fd}(g)\supseteq\operatorname{fd}(fg)$, but note that the inclusion may be strict. In any case, all divisors in $\operatorname{Int}(D)$ of an image-primitive polynomial $f\in\operatorname{Int}(D)$ are image-primitive.
@@ 4.1. No prime elements
We conclude from Fact 4.3 that every non-constant polynomial in $\mathbb{Z}[x]$ has zeros modulo infinitely many primes.
@@
Now consider $G\in\operatorname{Int}(\mathbb{Z})$ non-constant, $G=\frac{g}{d}$ with $g(x)\in\mathbb{Z}[x]$ and $d\in\mathbb{Z}$. Let $p\in\mathbb{Z}$ be prime such that $g$ has a zero modulo $p$ but the polynomial function induced by $g$ is not constant zero modulo $p$. (Such a prime $p$ exists because $g$ has a zero modulo infinitely many primes, but only finitely many primes divide the fixed divisor of $g$.)
[…]
Then, both $\frac{h(x)G(x)}{p}$ and $\frac{(G(x)+p)h(x)}{p}$ are integer-valued. This means that $G(x)$ divides $\frac{(G(x)+p)h(x)G(x)}{p}$, but $G(x)$ divides neither $G(x)+p$ (not even in $\mathbb{Q}[x]$) nor $\frac{h(x)G(x)}{p}$ (because $\frac{h(x)}{p}$ is not integer-valued).
@@
Since $\mathbb{Z}$ is a d-ring, it follows by Fact 4.5 that $\mathcal{O}_{K}$, too, is a d-ring.